There are two carotid arteries that feed blood to the brain, one on each side of the neck and head. One patient's carotid arteries are each 11.2 cm long and have an inside diameter of 5.2 mm . Near the middle of the left artery, however, is a 2.0-cm -long stenosis, a section of the artery with a smaller diameter of 3.5 mm .For the same blood flow rate, what is the ratio of the pressure drop along the patient's left carotid artery to the drop along his right artery?
Answer #1

Velocity is proportional to area and so proportional to the square of the diameter, and so the velocity on the left side is
V = v*(5.3/3.5)² = 2.293v


Bernoulli: p₀ + ρgh₀ + ½ρv₀² = p₁ + ρgh₁ + ½ρv₁²
so at the left side middle we have
p₀ - p₁ = ρg(h₁-h₀) + ½ρ(v₁² - v₀²) = ρ[g(h₁-h₀) + ½(v₁² - v₀²)]
p₀ - p₁ = ρ[9.8m/s²*0.112m/2 + ½(2.293² - 1)*v₀²
p₀ - p₁ = ρ[0.55m²/s² + 2.13v₀²]

and for the right side
p₀ - p₁ = ρg(h₁-h₀) + ½ρ(v₁² - v₀²) = ρg(h₁-h₀)
p₀ - p₁ = ρ*9.8m/s²*0.112m/2
p₀ - p₁ = ρ*0.55m²/s²

Well, I don't know how to find a ratio of these two values without knowing the velocity of the blood in the "normal" section.
If v₀ = 5 cm/s then the ratio is (0.55 + 2.13*.05²) / 0.55 = 1.01
But if v₀ = 20 cm/s, then the ratio is (0.55 + 2.13*.20²) / 0.55 = 1.16

given v₀ somewhere?

OR, if we ignore the effect of the height difference, the pressure drop is proportional to the square of the velocity, and so the left side pressure
P = p*(2.293v / v)² = 5.258p
so
P / p ≈ 5.26